# Deep learning modelling of manufacturing and build variations on multi-stage axial compressors aerodynamics

### v3

@@ Abstract
Application of deep learning methods to physical simulations such as CFD (Computational Fluid Dynamics) for turbomachinery applications, have been so far of limited industrial relevance. This paper demonstrates the development and application of a deep learning framework for real-time predictions of the impact of manufacturing and build variations, such as tip clearance and surface roughness, on the flow field and aerodynamic performance of multi-stage axial compressors in gas turbines. The associated scatter in compressor efficiency is known to have a significant impact on the corresponding overall performance and $CO_{2}$ emissions of the gas turbine, therefore posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance and potentially reduce requirements for expensive physical tests.
@@ 1 Introduction
Gas turbine manufacturers have gathered vast amount of operational data over the years, enabling them to establish best practices and design guidelines for their manufacturing and build processes. These guidelines, often consider a trade-off between product cost and engine performance, with tolerance ranges specified accordingly. However, the impact on performance used to define these ranges is typically based on previous experience and simplified correlations. Significant advancements in accuracy and computational cost, have also made CFD analyses an integral part of industrial design processes of turbomachinery components. Despite these advances, analytical models are generally limited to simplified scenarios (not considering real-world effects). Build-specific CFD models have been demonstrated to provide excellent agreement with test data Bruni2024CFD , but are not traditionally used as part of the manufacturing and build process as a day-to-day occurrence, due to the associated computational cost and requirement for specialized engineers to carry out the analyses. The proposed deep learning framework aims to model the effect of manufacturing and build variations on engine performance, achieving similar accuracies to standard CFD solvers, in a significantly shorter timescale. The focus of the current work is on tip clearance and surface roughness variations, two of the main sources of performance variability, but the framework is readily generalisable to other manufacturing variations. This is incorporated into the manufacturing and build process, providing instantaneous feedback, which can potentially be used to reduce the requirements for expensive physical testing when clearances go outside the acceptance limits.
The challenge: The impact of manufacturing and build variations on the overall performance of gas turbines is known to be significant Montomoli2018 Wang2020 , with the axial compressor, as shown in Figure 1, being a large contributor. For a given design, compressor efficiency is affected by a combination of in-tolerance variations, which can lead to a substantial increase in $CO_{2}$ emissions compared to the design target. The aerodynamic performance of multi-stage axial compressors is predicted as standard practice using CFD (Computational Fluid Dynamics). However, these calculations are traditionally performed only for a simplified ”as designed” geometry, that differs from the specific build of each engine. Therefore, accurately predicting the performance of each engine build by modelling the impact of manufacturing and build variations on engine performance within a short timescale is of great industrial and environmental importance.
Our contribution: This work demonstrates the application of a novel deep learning framework previously developed by the authors Bruni2023 for predicting the impact of manufacturing and build variations on the flow field and overall performance of a multi-stage axial compressors in real-time. Our proposed methodology achieves accuracy levels similar to those of a standard CFD solvers, within a significantly shorter timescale. This work extends the architecture first introduced in previous work by the authors Bruni2023 , from a proof of concept to a significantly more complex and computationally expensive application of great industrial relevance, demonstrating the aim to provide a generalisable framework. This is then proven to be readily integrated into the manufacturing and build process, providing the capability to model each engine build. In addition to avoiding the computational costs, long lead times, and specialized personnel requirements of traditional CFD analyses, it can be used to potentially reduce expensive physical tests.
@@
1.1 Tip Clearance variation background
Several manufacturing and build variations affect the overall performance of axial compressors, including tip clearance variations which are one of the main contributors Montomoli2018 Wang2020 . The term tip clearance refers to the radial distance between the stationary and rotating components within the compressor, such as the rotor blades and casing, or stator blades and rotor hub. Generally, the tip clearance must be large enough to prevent rubbing during operation, but small enough to avoid excessive aerodynamic losses. Axial compressors are designed to operate at a defined tip clearance, which can vary across stages. During manufacturing, the ”cold clearance” is specified and measured as part of the build process in stationary conditions. However, during operation, the ”hot clearance” can be significantly smaller due to thermal expansion and centrifugal effects. The cold clearance values have a target value and tolerance range, set by gas turbine manufacturers, based on trade-offs between undesirable rubbing and minimizing tip leakage losses. For a given ”cold clearance”, the corresponding ”hot clearance” can be calculated using either analytical models or operational experience. The ”hot clearance” is the actual clearance found during operation, and therefore is used in the CFD and machine learning models presented in this paper. The aerodynamic mechanisms driving tip leakage flows have been extensively researched over the years Storer1991 and can be modeled analytically with varying degrees of fidelity. Steady-state single passage CFD calculations are often used to model the effects of tip clearance variations on overall performance Sakulkaew2013 , and are used as ground-truth for this study. Non-axisymmetric tip clearance variations are known to affect the overall performance and surge margin Suriyanarayanan2022 and can be relevant for in-service degradation. However, for manufacturing variations they can be neglected Bruni2023 and single passage models can be used for the analyses.
1.2 Surface roughness variation background
CFD analyses traditionally assume aerofoil and gas path surfaces to be aerodynamically smooth. This assumption is often not valid for production components, which have a non-negligible roughness due to coating selection or manufacturing methods. A trade-off can be identified between production cost and roughness, as smoother finishes tend to be progressively more expensive. Therefore, it is of paramount importance to be able to characterize the impact of different surface roughness levels on the aerodynamic performance. This is particularly relevant when comparing test data from different builds with coating from different suppliers, as they can potentially have significantly different roughness. Experimental work has shown that increased roughness on the aerofoil surface has a detrimental effect on efficiency. One possible way to describe the surface roughness is to consider the average $R_{a}$ for each blade row, which can be measured for a given set of blades. The $R_{a}$ can then be converted to an equivalent sand grain roughness to be used for CFD using the equation $k_{s}=5.2\cdot R_{a}$. More details on surface roughness modelling for CFD predictions is provided by the authors in previous work Bruni2024CFD .
1.3 Machine Learning applications to turbomachinery
In recent years, there has been a significant amount of research into the utilization of machine learning techniques for turbomachinery applications. These include predicting the aerodynamic performance due to tip clearance variations Krishnababu2021 and design modifications Pongetti2021 , as well as aeromechanic predictions for both forced response Bruni2022 and flutter He2022 . As often the output variables are predicted directly from the input variables with a ”black box” approach, these methodologies have varying degrees of accuracy and generalisability, depending on the complexity of the application. However, the ability to accurately predict variations in target variables for turbomachinery in a manner that is both reliable and generalisable, is contingent upon the accuracy of flow field predictions. It is only through precise flow field predictions that the overall performance figure of interest can be calculated through the application of relevant physical equations that govern the dynamics. The potential benefits of using machine learning approaches for predicting full flow fields, have been demonstrated in some simplified cases in literature for 2D aerofoils on both cartesian Thuerey2018 Thuerey2020 Ribeiro2020 and unstructured Kashefi2021 grids. Li et. al. Li2023 , using CNNs to predict the flow field downstream of a single fan row, which was then used for noise predictions applications. A similar approach was used also by Rao et. al. Rao2023 for the semantic segmentation of an aero-engine intake using U-Net and U-Net++ architectures. Graph neural networks Harsch2021 Pfaff2021 have also been applied to 3D turbomachinery, with a single row compressor Perrone2022 and turbine Li2022 , which suggested potential challenges related to scalability and practicality for industrial applications. For instance, the computational cost associated with training, and the size of the required model, make the approach impractical for multi-stage compressors. These applications typically use computational meshes in the order of magnitude of 10 million nodes. Furthermore, storing the CFD data required for training and future simulations would require several petabytes, for a typical engine manufacturer, which is not practical. Purely data-driven approaches typical of supervised learning require large amount of data, so that the network can learn the solution manifold. A recent review thuerey2021pbdl has highlighted the potential of other approaches to mitigate these issues such as Physics-Informed Neural Networks (PINNs), Reinforcement Learning (RL), and Generative Adversarial Networks (GANs) Chu2021 . However, none of those methods can currently provide accurate, generalisable and scalable solutions to complex 3D turbomachinery applications.
1.4 Novel solution to machine learning scalability
To address these challenges, we introduce a pre-processing step to extract only the relevant engineering data from the computational domain, such as blade-to-blade planes, blade surfaces, axial cuts, loading distributions, radial distributions, stage-wise performance, and overall performance. In most cases, these provide sufficient information to the aerodynamicist for design and analyses activities, while close examination of the 3D flow field of a CFD solution is generally only necessary in non-standard cases. Once the relevant engineering data has been extracted from the computational domain, it can be interpolated onto a simplified grid with a resolution that is deemed acceptable for the application in question. Only selected fundamental variables are predicted, while the derived ones can be computed using the appropriate physical equations, significantly reducing the computational cost without sacrificing accuracy, as only the necessary data required to calculate the target variables is retained. Targeting the flow field predictions and using the relevant physical equations to calculate the corresponding overall performance renders the methodology generalisable, while filtering only relevant parts of the CFD solution makes the methodology scalable to complex industrial applications. This was first introduced by the authors in a previous work for a single stage compressor application Bruni2023 , and is now extended to a multi-stage compressor application, with a computational domain one order of magnitude larger. This paper demonstrates that the framework developed in the previous paper can be translated from an academic test case, to a complex real-world application with significant engineering relevance.
2 Methodology
A modern multi-stage industrial axial compressor consisting of 10 stages, as shown in Figure 1, is considered in this work. Each stage is comprised by a row of rotor blades and a row of stator blades with a given blade count. The effect of tip clearance and surface roughness variations on each blade row are the focus of this study. Future work will include other manufacturing and build variations such as geometry variations, as well as different operating conditions.
2.1 Data Generation
@@
The ground-truth data used for training are the CFD results for the configuration of interest. This computational model consisted of a single passage model of the 10 stages, with mixing plane interfaces. The computational mesh was generated using Numeca AUTOGRID5, and the CFD solver used was Trace, with SST $K-\omega$ turbulence model. More details on the computational setup used have been provided in previous publications Bruni2022 . The development of build-specific CFD models for compressor aerodynamic predictions and its validation against engine test data was also presented in previous work by the authors, discussing the relative effects of various manufacturing and build variations Bruni2024CFD . The tip clearance and surface roughness were varied for each row (I1=IGV, R1=Rotor-1, S1=Stator-1, etc.) within out-of-tolerance conditions up to 50% larger and tighter than the drawing specifications, as shown in Figure 2. This range is significantly larger than the tolerance typically specified for gas turbine applications, and was selected to demonstrate the robustness of the methodology to out-of-tolerance cases. The input variable space was sampled using latin-hypercubic sampling. The dataset generated for this study comprises 400 CFD solutions, each of which was executed using automated and parallelized processes, taking 90 minutes on 72 CPUs for each case. Even if the computational cost required to generate the dataset for supervised training is potentially high, it should be noted that most of the data is already available from internal work of the industrial partner. It is envisioned that the majority of the data used for training of this framework will be based on available data, with additional computations carried out only when required, to extend the dataset for wider generalisability.
@@
The CFD results are processed to only extract six variables, namely, Total Pressure ($P_{t}$), Total Temperature ($T_{t}$), Axial Velocity ($V_{x}$), Tangential Velocity ($V_{t}$), Radial Velocity ($V_{r}$), and Density ($\rho$), at each of the 24 inter-row locations as shown in Figure 1. The flow condition at each mesh node is fully described by the six variables and can be used to calculate all other available flow variables in the CFD solution. While the CFD mesh for each simulation comprises over 15 million nodes, the post-processed domain contains less than 100 thousand nodes overall, yet it includes all the necessary information for the desired engineering assessment. This consists only of the 24 inter-row 2D planes with a $64\times 64$ grid. The data is stored as a tensor with a channel-first convention and a shape of $(6\times 24\times 64\times 64)$, where the dimensions represent the number of variables, number of axial locations, tangential nodes in the mesh, and radial nodes in the mesh, respectively. To reduce the computational cost of the CFD analysis, rotational periodicity can be assumed, and only a single passage needs to be considered instead of the full annulus, as standard process for steady state CFD analysis. However, when generating single passage meshes, the location of the periodic boundaries can vary depending on the meshing approach or geometry. As the methodology aims to be generalisable, the data is interpolated for a given passage at a fixed circumferential location defined in the CFD post-processing step, irrespective of the periodic boundary definitions in the mesh. This is appropriate as the geometry definitions considered have a specific clocking configuration, and all the manufacturing and build variations are imposed based on it. This allows to obtain 2D contours which are not limited to a given mesh or geometry, as typical in most surrogate models developed for optimisations. However, 2D contours are only used for detailed analyses, whereas mass-flow averaging CumpstyAveraging in the circumferential direction is more commonly used to obtain radial profiles of the variables of interest for most engineering applications. These radial profiles of each variable are then mass-flow averaged CumpstyAveraging to obtain a 1D average, which is used to calculate stage-wise performance of the variables of interest, such as pressure-ratio ($PR$), and polytropic efficiency ($\eta_{p}$). In addition to the aerodynamic performance of the stage, it is possible to calculate the overall performance of the compressor. Predicting the stage-wise performance in addition to the overall performance is critical for implementation of the model in the gas turbine build process, as it provides detailed information on which blade rows should be targeted by corrective measures.

### v4

@@ Abstract
Applications of deep learning to physical simulations such as Computational Fluid Dynamics have recently experienced a surge in interest, and their viability has been demonstrated in different domains. However, due to the highly complex, turbulent and three-dimensional flows, they have not yet been proven usable for turbomachinery applications. Multi-stage axial compressors for gas turbine applications represent a remarkably challenging case, due to the high-dimensionality of the regression of the flow-field from geometrical and operational variables. This paper demonstrates the development and application of a deep learning framework for predictions of the flow field and aerodynamic performance of multi-stage axial compressors. A physics-based dimensionality reduction unlocks the potential for flow-field predictions, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. Compared to traditional ”black-box” surrogate models, it provides explainability to the predictions of overall performance by identifying the corresponding aerodynamic drivers. This is applied to model the effect of manufacturing and build variations, as the associated performance scatter is known to have a significant impact on $CO_{2}$ emissions, therefore posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data.
@@ 1 Introduction
The highly complex, turbulent and three-dimensional flow-field of multi-stage axial compressors has traditionally been one the main challenges for the application of CFD (Computational Fluid Dynamics) in the turbomachinery industry. In this context, the flow field consists of the aerodynamic variables of interest (such as pressure, temperature, velocities etc.) predicted on a numerical discretizations of the domain (i.e. mesh or grid), solving the Reynolds-Averaged Navier Stokes equations (RANS). Advancements in modelling and numerical methods have progressively improved its accuracy and computational cost, making CFD analyses an integral part of industrial design processes and greatly improving the design capabilities of gas turbine manufacturers. This has led to an interest in the development of robust design strategies to account for real world effects in the design and analyses, which are known to affect the overall performance as well as the operability of the compressor. Gas turbine manufacturers have gathered vast amount of operational data over the years, enabling them to establish best practices and design guidelines for their manufacturing and build processes. These guidelines, often consider a trade-off between product cost and engine performance, with tolerance ranges specified accordingly. However, the impact on performance used to define these ranges is typically based on previous experience and simplified correlations. The identification of the impact of manufacturing and build uncertainties on the operational performance can be used to both drive improvements in the manufacturing technology, and inform the design strategies to produce robust designs desensitized to these variabilities.
The challenge: Characterizing the impact of all the relevant manufacturing and build variations which are expected to affect the overall performance of a given compressor design, can be achieved by engine build-specific CFD models. While having demonstrated to provide excellent agreement with test data [1], these models are not traditionally used as part of the manufacturing and build process as a day-to-day occurrence, due to the associated computational cost and requirement for specialized engineers to carry out the analyses.
Our contribution: This paper demonstrates the development and application of a deep learning framework previously introduced by the authors [2] for predictions of the flow field and aerodynamic performance of multi-stage axial compressors, demonstrating the scalability from a proof of concept to a significantly more computationally expensive application of great industrial relevance. A physics-based dimensionality reduction unlocks the potential to use supervised learning approaches for flow-field predictions, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. The proposed architecture also has a competitive advantage compared to traditional ”black-box” surrogate models, as it provides explainability to the predictions of variations in overall performance by identifying the corresponding aerodynamic drivers. This is applied to model the effect of manufacturing and build variations, as the associated scatter in efficiency is known to have a significant impact on the overall performance and $CO_{2}$ emissions of the gas turbine, therefore posing a challenge of great industrial and environmental relevance. The proposed architecture is proven to achieve an accuracy comparable to that of the CFD benchmark, in real-time, for an industrially relevant application. The deployed model, is readily integrated within the manufacturing and build process of gas turbines, thus providing the opportunity to analytically assess the impact on performance with actionable and explainable data. This unlocks the capability to model each engine build avoiding the high computational costs, long lead times, specialized personnel requirements of traditional CFD analyses, and reducing the requirement for expensive physical tests.
1.1 Effect of manufacturing and build variations on the aerodynamic performance
The impact of manufacturing and build variations on the overall performance of gas turbines is known to be significant [3] [4], with the axial compressor being a large contributor. For a given design, compressor efficiency is affected by a combination of in-tolerance variations, which can lead to a substantial increase in $CO_{2}$ emissions compared to the design target. Therefore, accurately predicting the performance of each engine build by modelling the impact of manufacturing and build variations on engine performance within a short timescale is of great industrial and environmental importance. The focus of the current work is on tip clearance and surface roughness variations, two of the main sources of performance variability, but the framework is readily generalisable to other manufacturing variations. A detailed review of the effect of the various manufacturing and build variations on the compressor aerodynamic performance, and the associated modelling strategies has been previously been provided by the authors [1]. In the following sections, a brief overview is included to provide the readers with the necessary background to interpret the results presented in this paper.
1.1.1 Tip Clearance variation background
The term tip clearance refers to the radial distance between the stationary and rotating components within the compressor, such as the rotor blades and casing, or stator blades and rotor hub. Generally, the tip clearance must be large enough to prevent rubbing during operation, but small enough to avoid excessive aerodynamic losses. Axial compressors are designed to operate at a defined tip clearance, which can vary across stages. During manufacturing, the ”cold clearance” is specified and measured as part of the build process in stationary conditions. However, during operation, the ”hot clearance” can be significantly smaller due to thermal expansion and centrifugal effects. The cold clearance values have a target value and tolerance range, set by gas turbine manufacturers, based on trade-offs between undesirable rubbing and minimizing tip leakage losses. For a given ”cold clearance”, the corresponding ”hot clearance” can be calculated using either analytical models or operational experience. The ”hot clearance” is the actual clearance found during operation, and therefore is used in the CFD and machine learning models presented in this paper. The aerodynamic mechanisms driving tip leakage flows have been extensively researched over the years [5] and can be modeled analytically with varying degrees of fidelity. Standard practice in both academia and industry, is to use steady-state single passage CFD calculations to model the effects of tip clearance variations on overall performance [6], which are used as ground-truth for this study. This assumes that circumferential variations in tip clearance can be neglected. Overall, non-axisymmetric tip clearance variations are known to affect the overall performance and surge margin [7] and are of relevance especially for in-service degradation. However, for manufacturing variations they can be neglected [2] and single passage models can be used for the analyses.
1.1.2 Surface roughness variation background
Traditionally, CFD analyses are performed assuming aerofoil and gas path surfaces to be aerodynamically smooth. This assumption is often not valid for production components, which are bound to have non-negligible roughness due to coating selection or manufacturing methods. A trade-off can be identified between production cost and roughness, as smoother finishes tend to be progressively more expensive. Therefore, it is of paramount importance to be able to characterize the impact of different surface roughness levels on the aerodynamic performance. This is particularly relevant when comparing engine test data from different builds with coating from different suppliers, as they can potentially have significantly different roughness. Experimental work has shown that increased roughness on the aerofoil surface has a detrimental effect on efficiency. One possible way to model analytically the effect of surface roughness is to consider the average $R_{a}$ for each blade row, which can be measured for a given set of blades. The $R_{a}$ can then be converted to an equivalent sand grain roughness that can be prescribed in the CFD model on the airfoils surfaces.
1.2 Machine Learning applications to turbomachinery
In recent years, there has been a significant amount of research into the utilization of machine learning techniques for turbomachinery applications. These include predicting the aerodynamic performance due to tip clearance variations [8] and design modifications [9], as well as aeromechanic predictions for both forced response [10] and flutter [11]. As often the output variables are predicted directly from the input variables with a ”black box” approach, these methodologies have varying degrees of accuracy and generalisability, depending on the complexity of the application. However, the ability to accurately predict variations in target variables for turbomachinery in a manner that is both reliable and generalisable, is contingent upon the accuracy of flow field predictions. It is only through precise flow field predictions that the overall performance figure of interest can be calculated through the application of relevant physical equations that govern the dynamics. The potential benefits of using machine learning approaches for predicting full flow fields, have been demonstrated in some simplified cases in literature for 2D aerofoils on both cartesian [12] [13] [14] and unstructured [15] grids. CNNs have been demonstrated to predict the flow field downstream of a single fan row [16], which was then used for noise predictions applications. A similar approach was used also by Rao et. al. [17] for the semantic segmentation of an aero-engine intake using U-Net and U-Net++ architectures. Graph neural networks [18] [19] have also been applied to 3D turbomachinery, with a single row compressor [20] and turbine [21], which suggested potential challenges related to scalability and practicality for industrial applications. For instance, the computational cost associated with training, and the size of the required model, make the approach impractical for multi-stage compressors. These applications typically use computational meshes in the order of magnitude of 10 million nodes. Furthermore, storing the CFD data required for training and future simulations would require several petabytes, for a typical engine manufacturer, which is not practical. Purely data-driven approaches typical of supervised learning require large amount of data, so that the network can learn the solution manifold. A recent review [22] has highlighted the potential of other approaches to mitigate these issues such as Physics-Informed Neural Networks (PINNs), Reinforcement Learning (RL), and Generative Adversarial Networks (GANs) [23]. However, none of those methods can currently provide accurate, generalisable and scalable solutions to complex 3D turbomachinery applications.
1.3 Physics-based dimensionality reduction
To address these challenges, we introduce a pre-processing step to extract only the relevant engineering data from the 3D computational domain, such as: blade-to-blade planes, blade surfaces, axial cuts, loading distributions, radial distributions, stage-wise performance, and overall performance. In most cases, these provide sufficient information to the aerodynamicist for design and analyses activities, while close examination of the 3D flow field of a CFD solution is generally only necessary in non-standard cases. Once the relevant engineering data has been extracted from the computational domain, it can be interpolated onto a simplified grid with a resolution that is deemed acceptable for the application in question. The regression problem is therefore reformulated, from predicting the relevant variables on an unstructured computational grid in the order of magnitude of $\sim 10^{7}$ nodes, to a structured grid in the order of magnitude of $\sim 10^{5}$ nodes. Only selected fundamental variables are predicted, while the derived ones can be computed using the appropriate physical equations, significantly reducing the computational cost without sacrificing accuracy, as only the necessary data required to calculate the target variables is retained. Targeting the flow field predictions and using the relevant physical equations to calculate the corresponding overall performance renders the methodology generalisable, while filtering only relevant parts of the CFD solution makes the methodology scalable to complex industrial applications. This physics-based dimensionality reduction, unlocks the potential to use supervised learning approaches for turbomachinery applications, as it re-formulates the regression problem from an un-structured to a structured one, as well as reducing the number of degrees of freedom. The authors first introduced this approach in a previous work for a single stage compressor application [2], which is now extended to a multi-stage compressor application, with a computational domain one order of magnitude larger. The framework developed in the previous paper is demonstrated to be translated from an academic test case, to a complex real-world application with significant engineering relevance. While the focus of the current work is on multi-stage axial compressors, the framework is directly applicable to any other turbomachinery application, such as axial turbines.
2 Methodology
A modern multi-stage industrial axial compressor consisting of 10 stages, as shown in Figure 1, is considered in this work. Each stage is comprised of a row of rotor blades and a row of stator blades with a given blade count. The effect of tip clearance and surface roughness variations on each blade row are the focus of this study. Future work will include other manufacturing and build variations such as geometry variations, as well as different operating conditions. The extension of the model to a wider range of input features does not require in principle any alteration to the architecture and the framework, which has been developed to scale to an arbitrary number of variables.
2.1 Data Generation
The ground-truth data used for training are the CFD results for the configuration of interest. This computational model consisted of 10 stages, with mixing plane interfaces. To reduce the computational cost of the CFD analysis, rotational periodicity is assumed, and only a single passage needs to be considered instead of the full annulus, as standard process for steady state CFD analysis. An overview of the computational domain is shown in Figure 1, showing both the geometry of the airfoils as well as the axial velocity contours at the mixing plane locations. The computational mesh was generated using Numeca AUTOGRID5, and the CFD solver used was Trace, with SST $K-\omega$ turbulence model. More details on the computational setup used have been provided in previous publications [10]. The development of build-specific CFD models for compressor aerodynamic predictions and its validation against engine test data was also presented in previous work by the authors, discussing the relative effects of various manufacturing and build variations [1]. The quantification of the accuracy of the CFD simulations in comparison to test data is outside of the scope of the current work. The aim of the proposed architecture is to provide predictions of the flow field and overall performance of the compressor with accuracy similar to that of the CFD model, but without the associated computational cost.
@@ 2.1 Data Generation
The tip clearance and surface roughness were varied for each row (I1=Inlet Guide Vane, R1=Rotor-1, S1=Stator-1, etc.) within out-of-tolerance conditions up to 50% larger and tighter than the drawing specifications, as shown in Figure 2. This range is significantly larger than the tolerance typically specified for gas turbine applications, and was selected to demonstrate the robustness of the methodology to out-of-tolerance cases. The input variable space was sampled using latin-hypercubic sampling. The dataset generated for this study comprises 400 CFD solutions, each of which was executed using automated and parallelized processes, taking 90 minutes on 72 CPUs for each case. Even if the computational cost required to generate the dataset for supervised training is potentially high, it should be noted that most of the data is already available from internal work of the industrial partner. It is envisioned that the majority of the data used for training of this framework will be based on available data, with additional computations carried out only when required, to extend the dataset for wider generalisability. The use of a supervised learning approach is appropriate for industrial applications, as typically gas turbine manufacturers have access to a significant historical database of CFD calculations from previous analytical and design activities. In case of new applications for which historical data is not available, it’s advisable to favour physics-informed models, as the computational cost associated with generating the data from scratch would render supervised models impractical.
@@
2.2 Data pre-processing
As discussed in subsection 1.3, the CFD results are pre-processed to reduce the dimensionality of the problem, without any loss of information. First, only six variables are extracted, namely, Total Pressure ($P_{t}$), Total Temperature ($T_{t}$), Axial Velocity ($V_{x}$), Tangential Velocity ($V_{t}$), Radial Velocity ($V_{r}$), and Density ($\rho$). The flow condition at each mesh node is fully described by the six variables and can be used to calculate all other relevant flow variables in the CFD solution. Second, only the 24 axial locations shown in Figure 1 are considered. Third, as the methodology aims to be generalisable, the data is interpolated for a given passage at a fixed circumferential location, irrespective of the periodic boundary definitions, number of points and discretization used in the mesh. As the location of the periodic boundaries and distribution of the nodes can vary depending on the meshing approach or geometry, this approach is not limited to a given mesh or geometry, as typical in most surrogate models developed for optimisations. While the CFD mesh for each simulation comprises over 15 million nodes, the new pre-processed domain contains less than 100 thousand nodes overall, yet it includes all the necessary information for the desired engineering assessment. The new 3D flow field now consists now of only 24 axial 2D planes with a $64\times 64$ grid. The data is stored as a tensor with a channel-first convention and a shape of $(6\times 24\times 64\times 64)$, where the dimensions represent the number of variables, number of axial locations, radial nodes in the mesh, and tangential nodes in the mesh, respectively. Generally, 2D contours are only used for detailed analyses, whereas mass-flow averaging [24] in the circumferential direction is more commonly used to obtain radial profiles of the variables of interest for most engineering applications. These radial profiles of each variable are then mass-flow averaged radially to obtain a 1D average, which is used to calculate stage-wise performance of the variables of interest, such as pressure-ratio ($PR$), and polytropic efficiency ($\eta_{p}$). In addition to the aerodynamic performance of the stage, it is possible to calculate the overall performance of the compressor. Predicting the stage-wise performance in addition to the overall performance is critical for implementation of the model in the gas turbine build process, as it provides actionable data on which blade rows should be targeted by corrective measures.